Four times what Jill's age was three years ago is twice what she will be in 5 years? How old is she?

Answer by ikleyn(52781)

.

4*(J-3) = 2*(J+5)


4J - 12 = 2J + 10


4J - 2J = 10 + 12


2J      = 22


 J      = 22/2 = 11.


ANSWER.  Jill is 11 years old.

Solved.